# On the Brun–Titchmarsh theorem. I

### v3

@@ 4.1. Linear sieve
$S(\mathcal{A},\mathcal{P},z)\leqslant\mathscr{X}V(z)\bigg{\{}F\bigg{(}\frac{\log D}{\log z}\bigg{)}+E^{+}\bigg{\}}+\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}^{+}(d)r(\mathcal{A},d)$
@@
$S(\mathcal{A},\mathcal{P},z)\geqslant\mathscr{X}V(z)\bigg{\{}f\bigg{(}\frac{\log D}{\log z}\bigg{)}-E^{-}\bigg{\}}-\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}^{-}(d)r(\mathcal{A},d)$
@@ 4.2. Sifting primes in $\pi(x,q;a)$
(4.5) $\pi^{*}(x;q,a)=\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ (n,P(w))=1\end{subarray}}\varPhi(n)-\sum_{w\leqslant p<z}\sum_{\begin{subarray}{c}pn\equiv a\,(\operatorname{mod}{q})\\ (n,P(p))=1\end{subarray}}\varPhi(pn)+O(1+q^{-1}x^{\frac{1}{2}})$
@@
Let $\varepsilon>0$ sufficiently small and $D<x^{1-\varepsilon}.$ For any $2\leqslant w<z<\sqrt{x},$ we have
@@ 5.1. Approaching by multiplicative characters
(5.4) $-β\displaystyle\lambda^{+}=\boldsymbol{\alpha}*\bbeta,$
where $\boldsymbol{\alpha}$ and $-β\bbeta$ are of level $M$, $N$, respectively, and
@@
Here $\phi$ is a fixed smooth function which dominates the indicator function of $[-1,1].$
@@
$\sum_{p\sim P}\sum_{\begin{subarray}{c}pn\equiv a\,(\operatorname{mod}{q})\\ (n,P(p))=1\end{subarray}}\varPhi(pn)$ $\geqslant\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\end{subarray}}\lambda_{d}^{-}A(pd)$
$=\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\\ \omega(d)\leqslant(\log x)^{1/3}\end{subarray}}\lambda_{d}^{-}A(pd)+\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(p)\\ \omega(d)>(\log x)^{1/3}\end{subarray}}\lambda_{d}^{-}A(pd)=\pi_{1}^{*}+\pi_{2}^{*},$
say. Recall that $A(d)$ is defined by (4.6). Hence
$\pi_{2}^{*}$ $\ll\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ \omega(d)>(\log x)^{1/3}\end{subarray}}|\lambda_{d}^{-}|\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ pd\mid n\end{subarray}}\varPhi(n)$
$\ll 2^{-(\log x)^{1/3}}\sum_{d\leqslant 2PD^{*}}2^{\omega(d)}\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ d\mid n\end{subarray}}\varPhi(n)$
$\ll 2^{-(\log x)^{1/3}}\sum_{\begin{subarray}{c}n\leqslant 2x\\ n\equiv a\,(\operatorname{mod}{q})\end{subarray}}3^{\omega(n)}.$
This is at most $O(x/(q\log^{3}x))$ by Shiu [Sh80, Theorem 2], provided that $D>x^{0.01}$ and $q<x^{0.99},$ in which case we have
$\sum_{p\sim P}\sum_{\begin{subarray}{c}pn\equiv a\,(\operatorname{mod}{q})\\ (n,P(p))=1\end{subarray}}\varPhi(pn)$ $\geqslant\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\end{subarray}}\lambda_{d}^{-}A(pd)$
$=\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\\ \omega(d)\leqslant(\log x)^{1/3}\end{subarray}}\lambda_{d}^{-}A(pd)+O\Big{(}\frac{x}{q(\log x)^{3}}\Big{)}.$
@@
[…]
@@
[…]
@@
[…]
@@ 6.1. Approaching by additive characters
$-β\displaystyle\lambda=\boldsymbol{\alpha}*\bbeta,$
where $\boldsymbol{\alpha}$ and $-β\bbeta$ are of level $M$ and $N$, respectively, and
@@ 6.2. Transforming exponential sums
×$\ \ \ \ \times W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}m}}{q}\Big{)}.$
@@ 6.4. Bounding $\varSigma^{*}$ as trilinear forms
Each of the above choices of $N$ fails to improve the work of Iwaniec [Iw82] when $\varpi=1/2.$ This is due to the fact that we have explored the bilinear or trilinear structures in $\varSigma^{*},$ which requires that $K>q^{4\varepsilon}.$ This is impossible by recalling that $K=M^{-1}q^{1+\varepsilon}$ and $M=x^{1-\varepsilon}q^{-1}.$ We will beat this barrier in a forthcoming work ([XZ25]), i.e., we are able to improve Iwaniec [Iw82] when $\varpi=1/2.$
@@
The proofs of Theorems 2.1 and 2.2 would follow the same line, except that Lemma 2.3 applies in the proof of Theorem 2.1, and we utilize convolution structures of the coefficients when proving Theorem 2.2, so that Kloostermania (Lemma 2.4) applies.
@@ 7.1. Preparations
$\mathscr{F}_{\xi}(M,N)$ =$=\{(m,n)\in[1,M]\times[1,N]:(m,n)=1\text{ and }\omega(m),\omega(n)\leqslant\xi\}.$
Then there exists a partition of $\mathscr{F}_{\xi}(M,N)$ into at most $(\log(MN))^{2\xi}$ subsets $\mathscr{F}^{*}_{\xi}(M,N)$ having the following property:
@@
(7.1) $-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $=\sum_{h\leqslant H}\mathop{\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{r\leqslant R}\sum_{s\leqslant S}}_{(mn,rs)=(ns,a)=1}\alpha_{m,r}\beta_{h,n,s}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{rs}\Big{)}.$
The proof of Theorem 2.2 would require the following estimate.
@@
×$\ \ \ \ \times\Big{(}1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}\Bigg{]}^{\frac{1}{2}}$
@@
(7.2) $-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $=\sum_{\mathscr{F}^{*}_{\xi}(N,S)}\mathcal{Q}_{a}^{*}\{\mathscr{F}^{*}_{\xi}(N,S)\}.$
By Cauchy–Schwarz inequality, we have
@@
×$\ \ \ \ \times\mathop{\sum\sum}_{\begin{subarray}{c}r,m\in\mathbf{Z}\\ (ar,n_{1}n_{2})=(am,s_{1}s_{2})=(r,m)=1\end{subarray}}W\Big{(}\frac{r}{R}\Big{)}W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{ah_{1}\overline{mn_{1}}}{rs_{1}}-\frac{ah_{2}\overline{mn_{2}}}{rs_{2}}\Big{)}.$
@@
×$\ \ \ \ \times\mathop{\sum\sum}_{\begin{subarray}{c}r,m\in\mathbf{Z}\\ (ar,n_{1}n_{2})=(am,s_{1}s_{2})=(r,m)=1\end{subarray}}W\Big{(}\frac{r}{R}\Big{)}W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{ak\overline{mn_{1}n_{2}}}{rs_{1}s_{2}}\Big{)}.$
@@
$\Delta$ $=|a|^{\vartheta}\sqrt{MRS^{2}N^{2}}\frac{(1+\frac{MRS^{2}N^{2}}{|a|HNS})^{\vartheta}}{1+(\frac{|a|HNS}{MRS^{2}N^{2}})^{2}}\Big{(}1+\frac{|a|HNS}{MRS^{2}N^{2}}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}$
×$\ \ \ \times\Big{(}1+\frac{|a|HNS}{MRS^{2}N^{2}}+\frac{HNS}{N^{2}S^{2}}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}$
$=\sqrt{MN^{2}RS^{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}$
@@
$\mathcal{Q}_{1}$ $\ll\|\boldsymbol{\alpha}\|^{2}HNS((HNS)^{\frac{1}{2}}+NS)\Bigg{[}NS(MR)^{\frac{1}{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}$
×$\ \ \ \ \times\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}\Bigg{]}(HMNRS)^{\varepsilon},$
@@ 7.2. Initial reductions
$=\frac{1}{n_{2}}\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{h\in\mathbf{Z}}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta n_{1}}}{n_{2}}\Big{)},$
@@
$\widetilde{F}(\lambda;n_{1},n_{2},r)=\int_{\mathbf{R}}F\Big{(}y,\frac{n_{1}y-qr}{n_{2}}\Big{)}\mathrm{e}(-\lambda y)\mathrm{d}y.$
@@
×$\ \ \ \ \times\sum_{h\in\mathbf{Z}}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta n_{1}}}{n_{2}}\Big{)}.$
@@
(7.6) $=\frac{\varphi(q)}{q}\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d\\ (n_{1}n_{2},q)=(n_{1},n_{2})=1\end{subarray}}\frac{\alpha_{dn_{1}}\beta_{dn_{2}}}{n_{2}}\sum_{0\leqslant|r|\leqslant R/d}\widetilde{F}(0;n_{1},n_{2},r).$
@@
×$\ \ \ \ \times\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta n_{1}}}{n_{2}}\Big{)}.$
Denote by $\mathcal{R}^{*}_{1}$ the contribution to $\mathcal{R}^{*}$ from the term with $r=0$. Note that
$\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},0\Big{)}$ $\ll\eta L^{-1+\varepsilon}\delta n_{2}$
by (7.5). Hence
$\mathcal{R}^{*}_{1}\ll\eta(NL)^{1+\varepsilon},$
@@
(7.9) $\begin{split}\mathcal{R}^{*}_{2}&=\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\sum_{1\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d\\ (n_{1}n_{2},q)=(n_{1},n_{2})=1\end{subarray}}\frac{\alpha_{dn_{1}}\beta_{dn_{2}}}{n_{2}}\\ &\ \ \ \ \times\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta n_{1}}}{n_{2}}\Big{)}\end{split}$

### v4

@@ 4.1. Linear sieve
$S(\mathcal{A},\mathcal{P},z)\leqslant\mathscr{X}V(z)\Big\{F\Big(\frac{\log D}{\log z}\Big)+E^{+}\Big\}+\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}^{+}(d)r(\mathcal{A},d)$
@@
$S(\mathcal{A},\mathcal{P},z)\geqslant\mathscr{X}V(z)\Big\{f\Big(\frac{\log D}{\log z}\Big)-E^{-}\Big\}-\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}^{-}(d)r(\mathcal{A},d)$
@@ 4.2. Sifting primes in $\pi(x;q,a)$
(4.5) $\pi^{*}(x;q,a)\leqslant\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ (n,P(w))=1\end{subarray}}\varPhi(n)-\sum_{w\leqslant p<z}\sum_{\begin{subarray}{c}pn\equiv a\,(\operatorname{mod}{q})\\ (n,P(p))=1\end{subarray}}\varPhi(pn)+O(1+q^{-1}x^{\frac{1}{2}})$
@@
Let $\varepsilon>0$ sufficiently small and $D,D_{p}<x^{1-\varepsilon}.$ For any $2\leqslant w<z<\sqrt{x},$ we have
@@ 5.1. Approaching by multiplicative characters
[…]
@@
Here $\phi$ is a fixed smooth function which dominates the indicator function of [1,1].[-1,1].
@@
$\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\end{subarray}}\lambda_{d}^{-}r(pd)=\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\\ \omega(d)\leqslant(\log x)^{1/3}\end{subarray}}\lambda_{d}^{-}r(pd)+\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\\ \omega(d)>(\log x)^{1/3}\end{subarray}}\lambda_{d}^{-}r(pd)=\pi_{1}^{*}+\pi_{2}^{*}.$
Recall that $r(d)$ is defined by (4.7). Hence
$\pi_{2}^{*}$ $\ll\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ \omega(d)>(\log x)^{1/3}\end{subarray}}|\lambda_{d}^{-}|\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ pd\mid n\end{subarray}}\varPhi(n)+\frac{x}{q}\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ \omega(d)>(\log x)^{1/3}\end{subarray}}\frac{|\lambda_{d}^{-}|}{pd}$
$\ll 2^{-(\log x)^{1/3}}\sum_{d\leqslant 2PD^{*}}2^{\omega(d)}\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ d\mid n\end{subarray}}\varPhi(n)+\frac{2^{-(\log x)^{1/3}}x}{q}\sum_{d\leqslant 2PD^{*}}\frac{2^{\omega(d)}}{d}$
$\ll 2^{-(\log x)^{1/3}}\sum_{\begin{subarray}{c}n\leqslant 2x\\ n\equiv a\,(\operatorname{mod}{q})\end{subarray}}\tau^{2}(n)+\frac{2^{-(\log x)^{1/3}}x}{q}\sum_{d\leqslant 2PD^{*}}\frac{2^{\omega(d)}}{d}.$
Recalling that $P<z<\sqrt{x}$ and $D^{*}<x^{1-\varepsilon}$, the last expression is at most $O(x/(q\log^{2024}x))$ by Shiu [Sh80, Theorem 2], provided that $q<x^{0.99}.$ Thus we have
$\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\end{subarray}}\lambda_{d}^{-}r(pd)=\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\\ \omega(d)\leqslant(\log x)^{1/3}\end{subarray}}\lambda_{d}^{-}r(pd)+O\Big(\frac{x}{q(\log x)^{2024}}\Big).$
@@
Hence it remains to prove that for all $\in\{+,-\}$ and $1\leqslant k\leqslant 4,$
@@
(III) Treatments to $S_{3}^{\pm}$ and $T_{3}^{\pm}$: The treatments are similar to (II) by interchanging $M$ and $N$. Hence (5.11) with $\in\{+,-\}$ and $k=3$ holds as long as
@@
Hence (5.11) with $\in\{+,-\}$ and $k=4$ holds as long as $q\leqslant x^{\frac{9}{20}-\varepsilon}$.
@@ 6.1. Approaching by additive characters
[…]
@@ 6.2. Transforming exponential sums
$\ \ \ \ \times W\Big(\frac{m}{M}\Big)\mathrm{e}\Big(\frac{a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}m}}{q}\Big).$
@@ 6.4. Bounding $\varSigma^{*}$ as trilinear forms
Each of the above choices of $N$ fails to improve the work of Iwaniec [Iw82] when $\varpi=1/2.$ This is due to the fact that we have explored the bilinear or trilinear structures in $\varSigma^{*},$ which requires that $K>q^{4\varepsilon}.$ This is impossible by recalling that $K=M^{-1}q^{1+\varepsilon}$ and $M=x^{1-\varepsilon}q^{-1}.$ We will beat this barrier in the second paper of this series ([XZ25]), i.e., we are able to improve the work of Iwaniec [Iw82] when $\varpi=1/2.$
@@
The proofs of Theorems 2.1 and 2.2 will follow the same lines, except that Lemma 2.3 applies in the proof of Theorem 2.1, and we utilize convolution structures of the coefficients when proving Theorem 2.2, so that Kloostermania (Lemma 2.4) applies.
@@ 7.1. Preparations
$\mathscr{F}_{\xi}(M,N)$ $=\{(m,n)\in[1,M]\times[1,N]:(m,n)=1\text{ and }\omega(m),\omega(n)\leqslant\xi\}.$
Then there exists a partition of $\mathscr{F}_{\xi}(M,N)$ into at most $(2\log(3MN))^{\xi^{2}}$ subsets $\mathscr{F}^{*}_{\xi}(M,N)$ having the following property:
@@
[…]
The proof of Theorem 2.2 will require the following estimate.
@@
$\ \ \ \ \times\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{\frac{1}{2}}}\Big(1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big)^{\frac{1}{2}}\Big(1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big)^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}\Bigg]^{\frac{1}{2}}$
@@
[…]
Fix $W$ as a smooth function dominating the indicator function of $[1,2].$ By Cauchy–Schwarz inequality, we have
@@
$\ \ \ \ \times\mathop{\sum\sum}_{\begin{subarray}{c}r,m\in\mathbf{Z}\\ (ar,n_{1}n_{2})=(am,s_{1}s_{2})=(r,m)=1\end{subarray}}W\Big(\frac{r}{R}\Big)W\Big(\frac{m}{M}\Big)\mathrm{e}\Big(\frac{ah_{1}\overline{mn_{1}}}{rs_{1}}-\frac{ah_{2}\overline{mn_{2}}}{rs_{2}}\Big).$
@@
$\ \ \ \ \times\mathop{\sum\sum}_{\begin{subarray}{c}r,m\in\mathbf{Z}\\ (ar,n_{1}n_{2})=(am,s_{1}s_{2})=(r,m)=1\end{subarray}}W\Big(\frac{r}{R}\Big)W\Big(\frac{m}{M}\Big)\mathrm{e}\Big(\frac{ak\overline{mn_{1}n_{2}}}{rs_{1}s_{2}}\Big).$
@@
$\Delta$ $=|a|^{\vartheta}\sqrt{MRS^{2}N^{2}}\frac{(1+\frac{MRS^{2}N^{2}}{|a|HNS})^{\vartheta}}{1+(\frac{|a|HNS}{MRS^{2}N^{2}})^{\frac{1}{2}}}\Big(1+\frac{|a|HNS}{MRS^{2}N^{2}}+\frac{R}{MN^{2}}\Big)^{\frac{1}{2}}$
$\ \ \ \times\Big(1+\frac{|a|HNS}{MRS^{2}N^{2}}+\frac{HNS}{N^{2}S^{2}}\Big)^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}$
$=\sqrt{MN^{2}RS^{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{\frac{1}{2}}}\Big(1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big)^{\frac{1}{2}}\Big(1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big)^{\frac{1}{2}}$
@@
$\mathcal{Q}_{1}$ $\ll\|\boldsymbol{\alpha}\|^{2}HNS((HNS)^{\frac{1}{2}}+NS)\Bigg[NS(MR)^{\frac{1}{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{\frac{1}{2}}}\Big(1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big)^{\frac{1}{2}}$
$\ \ \ \ \times\Big(1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big)^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}\Bigg](HMNRS)^{\varepsilon},$
@@ 7.2. Initial reductions
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Denote by $\mathcal{R}^{*}_{1}$ the contribution to $\mathcal{R}^{*}$ from the term with $r=0$. Hence
[…]
Back to the situation before Poisson summation, we find
$\mathcal{R}_{1}^{*}$ $=\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d\\ (n_{1}n_{2},q)=(n_{1},n_{2})=1\end{subarray}}\alpha_{dn_{1}}\beta_{dn_{2}}\sum_{\begin{subarray}{c}l_{1}\equiv 0\,(\operatorname{mod}{n_{2}})\\ (l_{1},q)=1\end{subarray}}F\Big(l_{1},\frac{n_{1}l_{1}}{n_{2}}\Big)$
[…]
$=\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d\\ (n_{1}n_{2},q)=(n_{1},n_{2})=1\end{subarray}}\alpha_{dn_{1}}\beta_{dn_{2}}\sum_{(l_{1},q)=1}F(n_{2}l_{1},n_{1}l_{1})$
[…]
$\ll(NL)^{1+\varepsilon},$
@@
[…]